• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Hisham

    Log entry time 05:10:53 on October11,2010

    Entry number 334383

    keyword=Owl shift summary

    People:
    SL: H. Alabataineh
    TO: F. Itrad
    3rd: V. Bellini


    - Shift started during controlled access
    which lasted till 01:15
    -Got the beam back at 01:30
    -askedd for controlled access (escorted)
    to work on IOCs (Alex & Carlos)
    - A couple of white alarms
    -NMR lost locking at 02:37
    - at 04:12 gave up the beam for the rest of
    the shift due to magnert regulation.
    - 04:30asked for controlled access